OpenStudy (anonymous):

The sum of two numbers, x and y, is 12. The difference of x and two times y is 6. What are the values of x and y?

OpenStudy (mayankdevnani):

\[\large \bf 1)x+y=12\] \[2) \large \bf x-2y=6\]

OpenStudy (mayankdevnani):

can you solve simultaneous equations ? @zxcvbnm08765

OpenStudy (anonymous):

KINDA @mayankdevnani

OpenStudy (mayankdevnani):

From equation 2 ,you get \[\large \bf x=6+2y\]

OpenStudy (mayankdevnani):

right ??

OpenStudy (anonymous):

YEA

OpenStudy (mayankdevnani):

put this value of x into equation no.1,we get \[\large \bf 6+2y+y=12\] \[\large \bf 6+3y=12\]

OpenStudy (mayankdevnani):

can you figure out value of y ?

OpenStudy (anonymous):

THE ANSWER IS 2 RIGHT

OpenStudy (mayankdevnani):

correct

OpenStudy (mayankdevnani):

then put y=2,in \[\large \bf x=6+2y\]

OpenStudy (mayankdevnani):

then find out what is thee value of x ?

OpenStudy (mayankdevnani):

*the

OpenStudy (anonymous):

10

OpenStudy (mayankdevnani):

correct !!

OpenStudy (mayankdevnani):

so,we get our answer as :- x=10 and y=2

OpenStudy (mayankdevnani):

you can also give me MEDAL by clicking best response flashing right side to my comment !

OpenStudy (mayankdevnani):

@zxcvbnm08765

OpenStudy (mayankdevnani):

good !!